# ReflecTool: Towards Reflection-Aware Tool-Augmented Clinical Agents

### v1

@@ Abstract
Yusheng Liao♠,♢, Shuyang Jiang♣,♢, Yanfeng Wang♠,♢, Yu Wangthanks: Corresponding Author,♠,♢ Shanghai Jiao Tong University Fudan University Shanghai Artificial Intelligence Laboratory {liao20160907,wangyanfeng622,yuwangsjtu}@sjtu.edu.cn shuyangjiang23@m.fudan.edu.cn
@@ 1 Introduction
To address such shortcoming, numerous works developed more advanced clinical agents, which enable models to leverage complex information through specialized tools (Jin et al., 2024; Li et al., 2024a; Lin et al., 2024). For instance, EHRAgent (Shi et al., 2024) can access electronic health records (EHRs) via a code interface, and MMedAgent (Li et al., 2024a) can interpret medical images via several medical visual models (Li et al., 2024b; Ma et al., 2024). While these agents enhance LLMs’ ability to interact with various types of data, they remain limited to addressing specific clinical scenarios with a narrow range of tools, impeding their ability to interact with the diverse forms of information intrinsic to clinical environments (Hu et al., 2024; Lee et al., 2022; Adams et al., 2024). This lack of integration limits their effectiveness for further application in clinical scenarios.
@@ 2.1 Composition
The agents in clinical scenarios need to process the medical data in different formats, like medical images (Hu et al., 2024; Liu et al., 2021; Lau et al., 2018) and electronic health records (EHR) (Lee et al., 2022; Johnson et al., 2016; Pollard et al., 2018), to complete the analysis or diagnosis. However, previous works only focused on a simple scenario, with only limited types of tools to solve the problem (Shi et al., 2024; Li et al., 2024a; Tang et al., 2024). To approximate realistic clinical scenarios and evaluate the general capabilities of the agent in the medical field, we investigate existing public medical datasets and divide them according to the ability requirement of the agents. We built a benchmark term CAB, which contains five capacity dimensions and 18 tasks in total. Specifically, the definition of the five dimensions and the corresponding tasks are as follows:
Knowledge&Reasoning
MultiModal
Medical images [^1] The multimodal in this paper mainly indicates that medical images with multiple modalities. are a common form of information in clinical scenarios. Many diseases require a combination of image examination information to be accurately diagnosed. Here, we choose three datasets with 12 modalities and a wide range of task types. We chose SLAKE (Liu et al., 2021) and VQA-RAD (Lau et al., 2018) for the open-ended QA and OmniMedVQA (Hu et al., 2024) for the closed-ended QA.
Numerical Analysis
Data Understanding
Trustworthiness
@@ 2.2 Clinical Toolbox
Based on the proposed CAB, we develop a toolbox that contains 16 types of tools to enable agents to handle diverse tasks. For example, knowledge databases in the clinical toolbox enhance the medical knowledge of the agent, and calculators give the agent the ability to calculate indicators accurately. In order to allow the agent to solve problems more flexibly, we did not limit the types and number of tools when solving a specific type of task. Compared with medical agents that limit the tools to complete tasks (Li et al., 2024a; Jin et al., 2024), our method can give the agent better generalization and scalability. The details of the clinical toolbox are discussed in Appendix B.2.
@@ 3.3 Inference Stage
Models Know. MM. Num. Data. Trust. Total
Large Language Models
MedLlama3-8B 59.88 - 11.61 23.02 6.07 25.14
Qwen2-7B (Yang et al., 2024) 60.86 - 18.49 44.50 28.19 38.01
Llama3-8B (AI@Meta, 2024) 63.32 - 19.87 35.23 24.06 35.62
Llama3.1-8B (Dubey et al., 2024) 67.43 - 22.07 49.58 30.75 42.46
Qwen2-72B* (Yang et al., 2024) 72.90 - 31.07 50.61 40.45 48.76
Llama3.1-70B* (Dubey et al., 2024) 76.91 - 29.23 45.80 38.40 47.59
GPT-3.5-turbo (OpenAI, 2022) 63.64 - 19.18 24.26 18.17 31.31
MultiModal Large Language Models
MiniCPM-V-2.6 (Yao et al., 2024) 56.29 56.53 4.60 13.86 15.12 29.28
GPT-4o-mini (OpenAI, 2023) 73.65 48.47 29.61 50.05 57.91 51.94
Agent (Qwen2-7B)
COT (Wei et al., 2022) 58.91 - 19.98 36.17 44.68 39.94
ReAct (Yao et al., 2023) 62.03 49.47 24.05 29.24 53.87 43.73
CRITIC (Gou et al., 2024) 56.61 53.87 24.49 37.35 47.54 43.97
Reflexion (Shinn et al., 2023) 60.92 56.95 20.83 37.41 50.14 45.25
ReflecTool (Iterative Refinement, n=2) 63.79 60.83 21.97 51.65 48.60 49.37
ReflecTool (Candidates Selection, n=2) 62.81 61.91 26.78 52.20 41.72 49.08
Agent (Qwen2-72B*)
COT (Wei et al., 2022) 69.11 - 24.47 52.51 56.22 50.58
ReAct (Yao et al., 2023) 76.47 56.37 31.44 53.29 48.98 53.31
CRITIC (Gou et al., 2024) 74.01 54.96 30.92 55.15 46.69 52.35
Reflexion (Shinn et al., 2023) 76.79 60.95 31.99 58.37 53.75 56.37
ReflecTool (Iterative Refinement, n=2) 76.81 63.74 38.45 63.51 54.65 59.43
ReflecTool (Candidates Selection, n=2) 76.27 62.70 38.06 64.54 56.73 59.66
Reflect Type Reflective Memory Action-wise Reflection PubMedQA VQA-RAD EHRSQL MedMen MedHalt Avg.
@@
Table 3: Ablation results of Refinement and Selection verification methods. All the experiments are conducted on Qwen2-72B. The modules of the ReflecTool contain Reflective Memory and Action-wise Reflection.
@@ 5.3 Tool Distribution in Trajectory
To better investigate the impact of ReflecTool on tool selection, we visualize the tool distribution of different methods across various datasets in Figure 5. Since the task types in the Trustworthiness dimension overlap with the other four dimensions, we exclude Trustworthiness in the visualization of the tool distribution. The figure illustrates the advantages of ReflecTool. First, ReflecTool encourages the model to invoke a higher proportion of tools. Specifically, in the Knowledge & Reasoning dimension, ReAct tends to directly answer questions rather than utilize tools, whereas both variants of ReflecTool exhibit a higher rate of tool usage, leading to better task completion. Second, ReflecTool leads to more frequent use of similar types of tools. In the Knowledge dimension, ReAct and CRITIC tend to use only a limited number of tools from the same dimension, without attempting to leverage different tools to solve problems. This makes the model susceptible to the limitations of individual tools. In contrast, ReflecTool uses multiple tools of the same type to integrate information from diverse sources, thereby improving task performance. Finally, ReflecTool also demonstrates a higher proportion of invoking tools from different categories. This indicates that ReflecTool is more flexible in tool usage, being capable of experimenting with a broader range of tools to solve problems. It is worth noting that some tools cannot be used across dimensions. For instance, the ReAct in Data Understanding frequently calls the tool SQLCoder. However, despite the absence of database information in the Data Understanding dimension, SQLCoder can not obtain any information, and this is the incorrect usage of tools.
@@ B.1.2 MultiModal
OmniMedQA Hu et al. (2024) is a large-scale medical visual question-answering benchmark with 118,010 images and 127,995 question-answer pairs collected from 73 medical datasets covering 12 different imaging modalities and more than 20 anatomical regions. Here, we randomly sample the 1000 instances from the public part of the dataset and keep the data component proportions to avoid bias.
@@ B.1.5 Trustworthiness
Similar to EHR-Halt, LongHalt is also derived from the LongHealthQA following the operation described in Adams et al. (2024). We randomly sample multiple documentation except for the note containing the answer to the question. The model without the answer in the context can only refuse to answer the question.
@@ B.2 Clinical Toolbox
The description of the tool in the pre-built toolbox is shown in Table 4.
Appendix C Experiments Details
C.1 Implementations

### v2

@@ Abstract
Yusheng Liao,♠,♢, Shuyang Jiang♣,♢, Yanfeng Wang♠,♢, Yu Wangthanks: Corresponding Author,♠,♢ Shanghai Jiao Tong University Fudan University Shanghai Artificial Intelligence Laboratory {liao20160907,wangyanfeng622,yuwangsjtu}@sjtu.edu.cn shuyangjiang23@m.fudan.edu.cn
@@ 1 Introduction
To address such shortcoming, numerous works developed more advanced clinical agents, which enable models to leverage complex information through specialized tools (Jin et al., 2024; Li et al., 2024a; Lin et al., 2024). For instance, EHRAgent (Shi et al., 2024) can access electronic health records (EHRs) via a code interface, and MMedAgent (Li et al., 2024a) can interpret medical images via several medical visual models (Li et al., 2024b; Ma et al., 2024). While these agents enhance LLMs’ ability to interact with various types of data, they remain limited to addressing specific clinical scenarios with a narrow range of tools, impeding their ability to interact with the diverse forms of information intrinsic to clinical environments (Hu et al., 2024b; Lee et al., 2022; Adams et al., 2024). This lack of integration limits their effectiveness for further application in clinical scenarios.
@@ 2.1 Composition
The agents in clinical scenarios need to process the medical data in different formats, like medical images (Hu et al., 2024b; Liu et al., 2021; Lau et al., 2018) and electronic health records (EHR) (Lee et al., 2022; Johnson et al., 2016; Pollard et al., 2018), to complete the analysis or diagnosis. However, previous works only focused on a simple scenario, with only limited types of tools to solve the problem (Shi et al., 2024; Li et al., 2024a; Tang et al., 2024). To approximate realistic clinical scenarios and evaluate the general capabilities of the agent in the medical field, we investigate existing public medical datasets and divide them according to the ability requirement of the agents. We built a benchmark term CAB, which contains five capacity dimensions and 18 tasks in total. The details about the definition of the five dimensions and the corresponding tasks can be found in Appendix B.
@@ 2.2 Clinical Toolbox
Based on the proposed CAB, we develop a toolbox that contains 15 types of tools to enable agents to handle diverse tasks. For example, knowledge databases in the clinical toolbox enhance the medical knowledge of the agent, and calculators give the agent the ability to calculate indicators accurately. In order to allow the agent to solve problems more flexibly, we did not limit the types and number of tools when solving a specific type of task. Compared with medical agents that limit the tools to complete tasks (Li et al., 2024a; Jin et al., 2024), our method can give the agent better generalization and scalability. The details of the clinical toolbox are discussed in Appendix B.2.
@@
Models Know. MM. Num. Data. Trust. Total
Large Language Models
MedLlama3-8B 59.88 - 11.61 23.02 6.07 25.14
Qwen2-7B (Yang et al., 2024) 60.86 - 18.49 44.50 28.19 38.01
Llama3-8B (AI@Meta, 2024) 63.32 - 19.87 35.23 24.06 35.62
Llama3.1-8B (Dubey et al., 2024) 67.43 - 22.07 49.58 30.75 42.46
Qwen2-72B* (Yang et al., 2024) 72.90 - 31.07 50.61 40.45 48.76
Llama3.1-70B* (Dubey et al., 2024) 76.91 - 29.23 45.80 38.40 47.59
GPT-3.5-turbo (OpenAI, 2022) 63.64 - 19.18 24.26 18.17 31.31
MultiModal Large Language Models
MiniCPM-V-2.6 (Yao et al., 2024) 56.29 56.53 4.60 13.86 15.12 29.28
GPT-4o-mini (OpenAI, 2023) 73.65 48.47 29.61 50.05 57.91 51.94
Agent (Qwen2-7B)
COT (Wei et al., 2022) 58.91 - 19.98 36.17 44.68 39.94
ReAct (Yao et al., 2023) 62.03 49.47 24.05 29.24 53.87 43.73
CRITIC (Gou et al., 2024) 56.61 53.87 24.49 37.35 47.54 43.97
Reflexion (Shinn et al., 2023) 60.92 56.95 20.83 37.41 50.14 45.25
ReflecTool (Iterative Refinement, n=2) 163.79 160.83 121.97 151.65 148.60 149.37
ReflecTool (Candidates Selection, n=2) 162.81 161.91 126.78 152.20 41.72 149.08
Agent (Qwen2-72B*)
COT (Wei et al., 2022) 69.11 - 24.47 52.51 56.22 50.58
ReAct (Yao et al., 2023) 76.47 56.37 31.44 53.29 48.98 53.31
CRITIC (Gou et al., 2024) 74.01 54.96 30.92 55.15 46.69 52.35
Reflexion (Shinn et al., 2023) 76.79 60.95 31.99 58.37 53.75 56.37
ReflecTool (Iterative Refinement, n=2) 76.81 163.74 138.45 163.51 54.65 159.43
ReflecTool (Candidates Selection, n=2) 76.27 162.70 138.06 164.54 156.73 159.66
Table 2: Experimental results of four types of models on Clinical Agent Bench. The ‘COT’ method indicates the agent runs without the pre-built tools. ‘*’ indicates the models use 4-bit GPTQ quantization. ‘-’ means the model is not capable of solving such a task. The best results are Bold, while the second best results are underlining. $\dagger$ and $\ddagger$ indicate the p-value $<0.05$ and $<0.01$ comparing with the strongest baseline Reflexion, respectively.
@@ 3.3 Inference Stage
Reflect Type Reflective Memory Tool-wise Reflection PubMedQA VQA-RAD EHRSQL MedMen MedHalt Avg.
@@
Table 3: Ablation results of Refinement and Selection verification methods. All the experiments are conducted on Qwen2-72B. The modules of the ReflecTool contain Reflective Memory and Tool-wise Reflection.
@@ 5 Analysis
5.1 Tool Usage Error
Although we categorized tools based on the capability dimensions of the model (as shown in Table 6), using tools from other dimensions is not necessarily incorrect. For example, when addressing questions in the Knowledge dimension, it is reasonable to employ the NER model from the Data Understanding dimension to extract specialized terms, which can then be used to query a knowledge graph more effectively. However, there are still some cases of incorrect tool usage for specific tools: (1) invoking an MLLM without a medical image input, (2) using SQLCoder and DBManual without a database input, and (3) employing LongDocRAG without document file input. Based on these points, we analyze the behavior of different methods in selecting tools. As shown in Table 4, ReflecTool can significantly reduce the tool selection error on both step-level and task-level. The results support that the tool-wise reflection mechanism can improve agents’ ability to use appropriate domain tools. Besides, we also analyze the parameter error of tool usage in Appendix C.1.
Tool Selection Error $\downarrow$
Methods Step-Level Task-Level
ReAct (Yao et al., 2023) 1.44 4.03
Reflexion (Gou et al., 2024) 0.92 1.67
CRITIC (Shinn et al., 2023) 0.24 0.33
ReflecTool (Iterative Refinement) 0.06 0.15
ReflecTool (Candidate Selection) 0.02 0.08
Table 4: Tool Selection Error among agent-based methods. ‘Step-Level’ indicates the error rate of tool selection in each step, and ‘Task-Level’ indicates the proportion of whether the tool selection errors happen in the instance reasoning process.
Figure 3: Impact of the optimization step on refinement and selection methods. The average performance is calculated using the same datasets in the ablation study.
@@ 5.2 Effect of Optimization Step
Few-shot Methods Avg.
Standard ReAct (Yao et al., 2023) 57.50
Reflexion (Gou et al., 2024) 56.87
CRITIC (Gou et al., 2024) 50.78
ReflecTool (Iterative Refinement) 59.07
ReflecTool (Candidate Selection) 60.72
Long-Term Memory ReAct (Yao et al., 2023) 60.73
Reflexion (Gou et al., 2024) 62.20
CRITIC (Gou et al., 2024) 57.35
ReflecTool (Iterative Refinement) 59.76
ReflecTool (Candidate Selection) 63.31
Table 5: Impact of the long-term memory on agent-based methods. Note that the baseline methods only adopt the success trajectories from the long-term memory instead of the tool-wise experience. ‘Standard’ indicates the few-shot sample is a fixed successful trajectory.
@@ 5.3 Size of Verification methods
5.4 Impact of the Long-Term Memory
To better investigate the impact of long-term memory, we compare the proposed ReflecTool and other agent-based methods under different types of few-shot. As shown in Table 5, long-term memory can effectively improve the performance of all agent-based methods. However, the proposed methods, ReflecTool, still outperform all baselines with both types of few-shot samples. The results again show the effectiveness of the tool-wise reflection mechanism.
@@ 5.5 Tool Distribution in Trajectory
To better investigate the impact of ReflecTool on tool selection, we visualize the tool distribution of different methods across various datasets in Figure 5. Since the task types in the Trustworthiness dimension overlap with the other four dimensions, we exclude Trustworthiness in the visualization of the tool distribution. The figure illustrates the advantages of ReflecTool. First, ReflecTool encourages the model to invoke a higher proportion of tools. Specifically, in the Knowledge & Reasoning dimension, ReAct tends to directly answer questions rather than utilize tools, whereas both variants of ReflecTool exhibit a higher rate of tool usage, leading to better task completion. Second, ReflecTool leads to more frequent use of similar types of tools. In the Knowledge dimension, ReAct and CRITIC tend to use only a limited number of tools from the same dimension without attempting to leverage different tools to solve problems. This makes the model susceptible to the limitations of individual tools. In contrast, ReflecTool uses multiple tools of the same type to integrate information from diverse sources, thereby improving task performance. Finally, ReflecTool also demonstrates a higher proportion of invoking tools from different categories. This indicates that ReflecTool is more flexible in tool usage, being capable of experimenting with a broader range of tools to solve problems.
@@ B.1.2 MultiModal
Medical images [^3] The multimodal in this paper mainly indicates that medical images with multiple modalities. are a common form of information in clinical scenarios. Many diseases require a combination of image examination information to be accurately diagnosed. Here, we choose three datasets with 12 modalities and a wide range of task types. We chose SLAKE (Liu et al., 2021) and VQA-RAD (Lau et al., 2018) for the open-ended QA and OmniMedVQA (Hu et al., 2024b) for the closed-ended QA.
@@
OmniMedQA Hu et al. (2024b) is a large-scale medical visual question-answering benchmark with 118,010 images and 127,995 question-answer pairs collected from 73 medical datasets covering 12 different imaging modalities and more than 20 anatomical regions. Here, we randomly sample the 1000 instances from the public part of the dataset and keep the data component proportions to avoid bias.
@@ B.1.5 Trustworthiness
Similarly to EHR-Halt, LongHalt is also derived from LongHealthQA after the operation described in Adams et al. (2024). We randomly sample multiple documentation except for the note containing the answer to the question. The model without the answer in the context can only refuse to answer the question.
@@ B.2 Clinical Toolbox
In this section, we introduce the details of the implementation of each tool in the clinical toolbox proposed. The description of the tool in the pre-built toolbox is shown in Table 6.
B.2.1 Knowledge Tools
Google Search
We use the implementation of the python library googlesearch-python [^8] https://pypi.org/project/googlesearch-python/. To avoid the context becoming too lengthy due to retrieval results, we have only used the titles and abstracts of the first ten retrieved results.
Medrag
Following the implementation of Xiong et al. (2024), the knowledge base in our method is composed of PubMed [^9] https://pubmed.ncbi.nlm.nih.gov/, StatPearls [^10] https://www.statpearls.com/, and Textbooks (Jin et al., 2021). We adopt BM25111111https://github.com/facebookresearch/faiss (Robertson et al., 2009) as the retriever to search the information from the medical knowledge base.
DrugBank
We download the drug table from the website of DrugBankOnline [^12] https://go.drugbank.com/. The table consists of drugs and their relative information, including description, state, indication, dosages, and so on.
UMLS
For UMLS knowledge graph, we use the API provided by the National Institute of Health (NIH) [^13] https://www.nlm.nih.gov/research/umls/index.html.
B.2.2 MultiModal Tools
HuatuoGPT
HuatuoGPT is a medical MLLM for medical image understanding. We adopt HuatuoGPT-Vision-7B [^14] https://huggingface.co/FreedomIntelligence/HuatuoGPT-Vision-7B as the medical image information gather. HuatuoGPT will provide the answer to the question generated by the agents.
MedCaptioner
MedCaptioner (Xie et al., 2024) is a medical image captioner which can generate the caption without any query. Besides, it decomposes the image report into five parts, including the Modality Classification, Structure Detection, ROI Analysis, Leison Texture, and Local-global Relation. The structured report gives a comprehensive description and can help the agent to better understand the medical image.
MiniCPM
MiniCPM (Hu et al., 2024a) is the MLLM developed for general domains. Here, we choose general domain MLLM as a supplement to medical MLLM to provide more options for the agent and improve the robustness of image understanding.
B.2.3 Numerical Tools
Calculator
The calculator is used to receive mathematical expressions generated by the agent and execute the calculations using Python logic. This functionality is implemented with Python’s built-in eval function.
DBManual
For DBManual, we refer to the implementation of the SQL knowledge base in EHRAgent (Shi et al., 2024), providing a detailed description of each SQL database, including its tables and columns. This allows the model to utilize DBManual to understand the structure of each database and the semantics of its columns, thereby improving task performance.
SQLCoder
The role of SQL Coder is to convert the agent-generated intent into SQL queries and return the results retrieved from the database. In the implementation, the SQL Coder and the agent share the same model. Since the model’s conversion to SQL syntax is not always successful, the SQL Coder can make up to three attempts based on the error messages encountered.
B.2.4 Data Tools
Spacy
SpaCy is used to extract all medical-specific terms from the input paragraph. To enhance the efficiency of entity extraction, we employ the en_core_sci_sm [^15] https://allenai.github.io/scispacy/ model from SciSpaCy [^16] https://github.com/allenai/scispacy as the NER model.
LongDocRAG
LongDocRAG is utilized to divide multiple user-uploaded documents into chunks and perform retrieval using Retrieval-Augmented Generation (RAG), enabling the agent to handle long contexts. In the implementation, we employ Llama-Index [^17] https://www.llamaindex.ai/ to accomplish this operation. It is worth noting that Llama-Index can handle multimodal data; however, in this work, we limit its use to processing textual data only.
B.3 Further Discussion
B.3.1 Clinical Tool
For the coverage of tools in the proposed Clinical Toolbox, it is intractable to account for any existing medical tool due to the diversity and complexity of medical tasks. However, to ensure that our work can be broadly applicable across various clinical scenarios, we have made a conscious effort to include a diverse range of tools. For Knowledge & Reasoning dimension, where task types are relatively uniform, we constructed a set of heterogeneous knowledge tools, including free text (Medrag), knowledge graphs (UMLS), tables (DrugBank), and search engines (Google Search). For the MultiModal dimension, three On the other hand, dimensions with inherently diverse task types naturally feature a wider variety of tools. Furthermore, our proposed method, ReflecTool, imposes no restrictions on tool types, enabling it to generalize effectively to other tools. In terms of MultiModal dimensions, the three selected multimodal models exhibit distinct areas of specialization. HuatuoGPT is tailored to the medical domain, providing domain-specific capabilities. MiniCPM serves as a general-purpose model, supplementing the medical model to enhance the diversity and functionality of the tools. MedCaptioner, designed specifically for generating descriptions of medical images, operates independently of queries provided by agents, thereby offering a unique utility within the framework. For other dimensions defined by diversity, a variety of tools is naturally required to ensure the comprehensive diversity of tools within the Clinical Toolbox.
B.3.2 Clinical Scenarios
Following the discussion on the comprehensiveness of tools, this section examines the medical scenarios in which agents can be applied. The five proposed dimensions correspond to five distinct types of scenarios anticipated for agent deployment. The Knowledge & Reasoning dimension equips the model with the ability to leverage medical knowledge for tasks such as medical question-answering and reasoning-based decision-making in clinical diagnoses. The MultiModal dimension enables the model to process diverse types of medical images. For instance, the selected multimodal medical models, HuatuoGPT and MedCaption, support the analysis of all common imaging modalities, including X-rays, CT scans, and MRI, thereby ensuring robust coverage in medical image interpretation. The Numerical Analysis dimension allows the model to integrate with hospital Electronic Health Record (EHR) systems, enhancing its capacity to assist physicians in delivering medical services. This capability also improves the model’s proficiency in interpreting numerical data, enabling accurate evaluations of clinical indicators for normalcy. The Data Understanding dimension enhances the model’s ability to process medical text, such as patient reports and long-form records like longitudinal follow-ups. Lastly, the Trustworthiness dimension focuses on minimizing hallucinations during medical task execution, thereby improving the model’s reliability in clinical applications.
Appendix C Case Study
C.1 Parameter Error in Tool Usage
There are two types of errors in the process of an agent utilizing tools to solve problems. The first is tool selection errors, as discussed in Section 5.1, where the agent employs an inappropriate tool for the task. The second is parameter errors, which occur when the agent uses the correct tool but fails to provide appropriate parameters, leading to unexpected results. A common example is when an agent invokes a knowledge tool but generates an unsuitable query, causing the tool to retrieve knowledge with low relevance to the task, thereby failing to provide the agent with effective information. Figure 7 presents a case study to illustrate this phenomenon more clearly. It can be observed that while both agents invoked the same tool, Medrag, Reflexion-Qwen2-72B failed to obtain the correct result. In contrast, ReflecTool-Qwen2-72B, despite initially failing to retrieve relevant information, successfully generated the correct query after reflection, ultimately retrieving effective information. This demonstrates that ReflecTool not only enhances the model’s ability to select the appropriate domain-specific tools but also enables the agent to leverage the tool-wise reflection mechanism to provide better parameters when invoking tools, thereby maximizing their effectiveness.
Figure 7: Case study for parameter error in tool usage. The figure compares the reasoning trajectories of two agent-based methods for the same task.
Appendix D Experiments Details
D.1 Implementations